# Spectroscopy and complex-time correlations using minimally entangled typical thermal states

### v1

@@ IV.2 Analytic complex-time correlation functions
Figure 8: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the auxiliary state |$\ket{\phi_{i}^{k}}$ Eq. 64 for evaluating the analytical correlation function $\mathcal{A}(z)$. We consider the $16\times 4$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ and $J/J_{D}=0.63$ and compare results for temperatures (a) $T/J_{D}=0.1$ and (b) $T/J_{D}=0.2$. The average is computed from $R=100$ METTS samples and the shaded region shows the estimated standard deviation of the mean.
@@ IV.3 Stochastic analytic continuation
To retrieve the spectral function from $\mathcal{A}(z)$ we perform analytic continuation, for which various approaches have been proposed and tested in the context of quantum Monte Carlo simulations [74, 75, 76, 77]. Although mostly used for performing continuation from imaginary time correlation functions, these approaches are often also applicable for the analytic correlator $\mathcal{A}(z)$. Generally, inverse Laplacian approaches from imaginary time correlation functions lead to a biased result where the details of the efficiency are case-dependent. In most cases, dominant peaks capturing well-defined quasiparticles can be resolved in a satisfactory way. However, for a spectrum with contributions from both a low energy peak and a higher energy continuum, numerical continuation is hard to quantitatively resolve the information at high frequency. On the other hand, we propose the numerical analytic continuation to be more efficient via correlation functions defined on the complex time case, since real-time contributions of the correlator, even for short periods of time, carry information about the high-frequency spectrum.
@@
where $\alpha$ denotes the fictitious inverse temperate, and $\chi^{2}(A)$ is the $\chi^{2}$ that measures the deviation of correlation function based on any given $A(\omega)$ during stochastic sampling, with the physical one that is observed from METTS calculation. In the current case, fluctuations from both real and imaginary parts of the correlation function are included in the definition of $\chi^{2}$,
@@
where $\mathcal{A}^{i}(z_{k})$ is the correlation function measured from the $i$-th METTS configuration.
@@
We show results of analytic continuation at $T/J_{D}=0.1$ and $0.2$ for several complex angles in Figs. 10 and 11. First, the spectrum obtained via the sampling at $\alpha_{c}$ does not differ from the one via weighted averaging the data from high to low temperature. Clearly, the position of the lowest triplon mode at $T/J_{D}=0.1$ is constant with both perturbation theory and the one from hermitian correlation function (at $\theta\longrightarrow 0$ limit) as shown in Fig. 12. Meanwhile, the second triplon peak at a higher frequency, which is not clearly visible for the case of $\theta=\pi/2$, can be resolved better from complex angle calculations. On the other hand, we did not observe a significant $\theta$ dependence for the thermal broadened spectrum at $T/J_{D}=0.2$.
@@ IV.4 Hermitian complex-time correlation functions
To evaluate this correlation function we can proceed analogously to the conventional time-dependent correlation function algorithm in Alg. 1, and define the states,
@@
using Alg. 1 adjusted for the complex time evolution along the contour $z=|z|e^{-i\theta}$ shown in Fig. 7. A key difference to the dynamical METTS algorithm Alg. 1 is the fact that the time evolution operators $e^{-iHz}$ and $e^{iH\overline{z}}$ are not norm preserving. However, we could reinforce a normalization by setting,
@@
We show hermitian spectral functions $S^{z}_{\theta}(\bm{q},\omega)$ for the $16\times 4$ Shastry-Sutherland lattice at $J/J_{D}=0.63$ and temperatures $T/J_{D}=0.1,0.2$ in Fig. 12. Remarkably, for $T/J_{D}=0.1$ the position of the dominant peak at $\omega/J_{D}\approx 0.46$ is always obtained consistently even for sizable values of the complex angle $\theta$. Meanwhile, the height and width of the dominant peak are well captured as long as $\theta\leq\pi/16$. Moreover, even features like secondary peaks at higher temperatures are captured correctly for not too large values of the complex angle $\theta$. Similarly, the overall shape of the exact spectral function with $\theta=0$ at $T/J_{D}=0.2$ in Fig. 12(b) is generally well captured at larger values of $\theta$. However, the peak around the triplon gap for $\omega/J_{D}\approx 0.46$ is overestimated at finite angles of $\theta$.
Figure 13: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the states |$\ket{\tilde{v}_{i}(z)}$ as in Eq. 88 for computing the hermitian correlator Eq. 43. Data on the $16\times 6$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ at $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$. We compare results for temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b). For larger values of $\theta$ we observe convergence of $S_{\text{vN}}$ towards the ground state entanglement entropy.
@@ V Discussion and Perspectives
On the other hand, the definition of the hermitian correlator has the advantage of easy implementation and the imaginary time evolution dissipates the entanglement entropy dramatically. Meanwhile, the main structure of the real frequency spectrum is well captured even in the presence of a systemic deviation due to imaginary contributions. Only recently, this version of a complex time correlator has been studied [53] in the context of impurity models with MPS.
It is worth pointing out, that the idea of introducing an imaginary-time component to reduce entanglement is similar in spirit to the dissipation-assisted operator evolution proposed in Ref. [45]. The imaginary-time component acts as a natural ”dissipator”. One advantage of the present approach is that existing implementations of time-evolution algorithms, such as TDVP, can be readily employed to compute complex-time evolution.
The SAC method based on the complex plane correlation function has still leftover possibilities to be improved in the future: one promising suggestion is to perform continuation via combined information of correlators at several angles along the complex plane. Apart from the stochastic approach, an interesting direction is to combine the Nevanlinna analytic continuation [78] approach with the correlation functions obtained at certain METTS snapshots. On the other hand, a common technique for computing dynamical spectral functions at zero temperatures by using real-time evolution with MPS, is to perform linear prediction in order to extend the time-series of correlation functions [79, 39, 80, 81]. While it is reported that this works remarkably well for simple spectra with few poles, we have not been able to successfully apply this to the correlation functions computed at non-zero temperatures in this work. For generic METTS states we mostly find strong disagreement between the actual time-evolution and the linear prediction. This is likely due to the fact that the correlators $\mathcal{C}_{i}(t)$ from METTS in general have a more complex structure than a superposition of a few sinusoidal waves. Thus, all the data we present is taken from actual time evolution. Recently, however, a more advanced method for performing linear prediction in the context of spectral functions with increased accuracy has been proposed [81]. It would be interesting to find out whether this method would be beneficial in the present context to extend the METTS time series.
@@ Appendix A Projection error of TDVP
Figure 14: Imaginary time correlation function using TEBD and the 2-site TDVP approach, for $T/\Delta=0$ (a), $0.1$ (b) and $0.2$ (c). Two dashed lines, including $c_{2}e^{-\Delta\tau}$ ( $\Delta\approx 0.459$ ) and $c_{1}e^{-0.56\tau}$ are guides to the eye. The time evolution approach for measuring correlation function at $T=0$ is based on the ground state via DMRG calculation, where the ones at $T/J_{D}=0.1$ and $0.2$ are from averaging the METTS configurations.
One potential error of the TDVP method exists at the step of projecting the time-dependent Schrödinger equation onto the relevant MPS subspace [67]. This projection error is especially severe when time evolving an MPS with very low bond dimension. It turns out to be a serious issue for the Shastry-Sutherland model when the temperature is much lower than the first triplon gap $\Delta$. In this low energy case, the MPS during the time evolution step does not differ significantly from the ground state, which is known to be a product state of spin singlets living on the off-diagonal bonds. The projection error leads to a biased estimation of the spin gap, even with the 2-site TDVP approach. On the other hand, time-evolving block decimation (TEBD) suffers only from the Trotter error and the MPS truncation error.
As a test, we show a comparison between TDVP and TEBD methods in Fig. 14. At low temperatures, one expects the imaginary time correlation function to decay with an exponential factor based on the estimated spin gap ($\Delta\approx 0.459J_{D}$) from the perturbation theory. However, as can be seen from the plot, $G(\tau)$ obtained from the 2-site TDVP method deviates from this expectation with a much steeper slope, for $T/\Delta=0$ and $0.1$. On the other hand, the long-time correlation from TEBD resolves the expected spin gap. Meanwhile, for temperatures comparable with the spin gap, enough entanglement is built up within the time-evolved MPS and we do not observe a severe projection error in TDVP. This can be seen from the consistent data between the two methods at $T/\Delta=0.2$. For the three temperatures here, we choose the time step in TEBD as $d\tau=0.1$ such that the Trotter error is sufficiently small.

### v2

@@ IV.2 Analytic complex-time correlation functions
Figure 8: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the auxiliary state |$\ket{\phi_{i}^{k}}$ Eq. 60 for evaluating the analytical correlation function $\mathcal{A}(z)$. We consider the $16\times 4$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ and $J/J_{D}=0.63$ and compare results for temperatures (a) $T/J_{D}=0.1$ and (b) $T/J_{D}=0.2$. The average is computed from $R=100$ METTS states and the shaded region shows the estimated standard deviation of the mean.
The key quantity determining the efficiency of the algorithm is the bipartite von-Neumann entanglement entropy $S_{\text{vN}}$ as in Eq. 36 of the MPS |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$, shown in Figs. 8 and 9, respectively. The entanglement is considered for a bipartition at the center of the system. We evaluated the analytic correlator for the dynamical spin structure factor $S^{z}(\bm{q},\omega)$ as in Eq. 3 for the $16\times 4$ Shastry-Sutherland model at $J/J_{D}=0.63$ at temperatures $T/J_{D}=0.1$ and $T/J_{D}=0.2$. The auxiliary state |$\ket{\phi_{i}^{0}}=\ket{\sigma_{i}}$ is a classical product state such that the entropy in Fig. 8 starts from zero. Slower entanglement growth is observed for larger values of $\theta$ as expected. For $\theta=\pi/2$ we are performing fully imaginary time evolution, which yields the slowest entanglement growth. Meanwhile, fluctuations of $S_{\text{vN}}$ among the sampled states during METTS dynamics are rather small, as indicated by the shaded regions around the mean. The absolute values of $S_{\text{vN}}$ are comparable for both |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$ and grow with increasing temperature. We observe a significant reduction in entanglement entropy when comparing the values to the entanglement of the states $|v_{i}(t)\rangle$ Eq. 18 as shown in Fig. 6. While we observed maximal values of $S_{\text{vN}}\approx 5.3$ for $T/J_{D}=0.2$ for $|v_{i}(t)\rangle$ in Fig. 6, we only find maximal values of $S_{\text{vN}}\approx 2.3$ for $\theta=\pi/8$ in Figs. 8 and 9. Naturally, the chosen time horizon $\Omega$ is significantly shorter when evaluating the analytical correlator. However, this demonstrates that $\mathcal{A}(z)$ can be evaluated at significantly reduced cost at non-zero values of $\theta$. We want to point out that setting a value of $\theta$ determines the value of $\Omega$ due to the constraint $\Omega_{\tau}=\beta/2$ we impose for the analytical correlator. Hence, the maximal complex times $|z|$ differ for different values of $\theta$. In our simulations, the maximal computed entanglement entropy exhibits a plateau due to simulating at finite bond dimension.
@@ IV.3 Stochastic analytic continuation
To retrieve the spectral function from $\mathcal{A}(z)$ we perform analytic continuation, for which various approaches have been proposed and tested in the context of quantum Monte Carlo simulations [74, 75, 76, 77]. Although mostly used for performing continuation from imaginary time correlation functions, these approaches are often also applicable for the analytic correlator $\mathcal{A}(z)$. Generally, inverse Laplacian approaches from imaginary time correlation functions lead to a biased result where the details of the efficiency are case-dependent. In most cases, dominant peaks capturing well-defined quasiparticles can be resolved satisfactorily. However, for a spectrum with contributions from both a low energy peak and a higher energy continuum, numerical continuation faces difficulties in quantitatively resolving the information at high frequency. On the other hand, we propose the numerical analytic continuation to be more efficient via correlation functions defined on the complex time case. At first sight, this natural extension would improve the performance of conventional imaginary-time-based numerical analytic continuation approaches since Green’s functions that we use as input are closer to the desired real frequency angle. Most importantly, the imaginary time Green’s function is dominated by the low-energy spectrum of the system. Instead, the high-energy part of the spectrum contributes significantly to the real-time correlator. Hence analytic continuations in our general consideration, with real-time contributions of the correlator, even for short periods, may give an improvement in resolving the spectrum.
@@
where $\alpha$ denotes the fictitious temperature, and $\chi^{2}(A)$ is the $\chi^{2}$ that measures the deviation of correlation function based on any given $A(\omega)$ during stochastic sampling, with the physical one that is observed from METTS calculation. In the current case, fluctuations from both real and imaginary parts of the correlation function are included in the definition of $\chi^{2}$,
@@
where $\mathcal{A}^{i}(z_{k})$ is the correlation function measured from the $i$-th METTS state.
@@
We show results of analytic continuation at $T/J_{D}=0.1$ and $0.2$ for several complex angles in Figs. 10 and 11. First, the spectrum obtained via the sampling at $\alpha_{c}$ does not differ from the one via weighted averaging the data from high to low temperature. Clearly, the position of the lowest triplon mode at $T/J_{D}=0.1$ is constant with both perturbation theory and the one from the Hermitian correlation function (at $\theta\longrightarrow 0$ limit) as shown in Fig. 12. Meanwhile, the second triplon peak at a higher frequency, which is not clearly visible for the case of $\theta=\pi/2$, can be resolved better from complex angle calculations. On the other hand, we did not observe a significant $\theta$ dependence for the thermal broadened spectrum at $T/J_{D}=0.2$.
The optimal angle for calculating the analytic correlation function is a question of balancing the increased computational cost at a smaller complex angle and the higher uncertainty in analytical continuation at a larger one. Certainly, among the numerical cost-accessible regions in the complex time plane, one may push the calculation to the smallest possible complex angle to acquire the most accurate spectral function.
@@ IV.4 Hermitian complex-time correlation functions
To evaluate this correlation function we can proceed analogously to the conventional time-dependent correlation function algorithm in Fig. 2, and define the states,
@@
using Fig. 2 adjusted for the complex time evolution along the contour $z=|z|e^{-i\theta}$ shown in Fig. 7. A key difference to the dynamical METTS algorithm Fig. 2 is the fact that the time evolution operators $e^{-iHz}$ and $e^{iH\overline{z}}$ are not norm preserving. However, we could reinforce a normalization by setting,
@@
We show Hermitian spectral functions $S^{z}_{\theta}(\bm{q},\omega)$ for the $16\times 4$ Shastry-Sutherland lattice at $J/J_{D}=0.63$ and temperatures $T/J_{D}=0.1,0.2$ in Fig. 12. Remarkably, for $T/J_{D}=0.1$ the position of the dominant peak at $\omega/J_{D}\approx 0.458$ is always obtained consistently even for sizable values of the complex angle $\theta$. Meanwhile, the height and width of the dominant peak are well captured as long as $\theta\leq\pi/16$. Moreover, even features like secondary peaks at higher temperatures are captured correctly for not too large values of the complex angle $\theta$. Similarly, the overall shape of the exact spectral function with $\theta=0$ at $T/J_{D}=0.2$ in Fig. 12(b) is generally well captured at larger values of $\theta$. However, the peak around the triplon gap for $\omega/J_{D}\approx 0.458$ is overestimated at finite angles of $\theta$.
Figure 13: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the states |$\ket{\tilde{v}_{i}(z)}$ as in Eq. 84 for computing the Hermitian correlator Eq. 44. Data on the $16\times 6$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ at $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$. We compare results for temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b). For larger values of $\theta$ we observe convergence of $S_{\text{vN}}$ towards the ground-state entanglement entropy.
@@ V Discussion and Perspectives
On the other hand, the definition of the Hermitian correlator has the advantage of easy implementation and the imaginary time evolution dissipates the entanglement entropy dramatically. Meanwhile, the main structure of the real-frequency spectrum is well captured even in the presence of a systemic deviation due to imaginary contributions. Only recently, this version of a complex time correlator has been studied [53] in the context of impurity models with MPS. Moreover, Ref. [53] provides a method of extrapolating correlation functions from the complex plane to the real axis which demonstrably is capable of achieving high accuracy.
It is worth pointing out that the idea of introducing an imaginary-time component to reduce entanglement is similar in spirit to the dissipation-assisted operator evolution proposed in Ref. [45]. The imaginary-time component acts as a natural “dissipator”. One advantage of the present approach is that existing implementations of time-evolution algorithms, such as TDVP, can be readily employed to compute complex time evolution.
The SAC method based on the complex plane correlation function has still leftover possibilities to be improved in the future: one promising suggestion is to perform continuation via combined information of correlators at several angles along the complex plane. Apart from the stochastic approach, an interesting direction is to combine the Nevanlinna analytic continuation [78] approach with the correlation functions obtained at certain METTS states. On the other hand, a common technique for computing dynamical spectral functions at zero temperatures by using real-time evolution with MPS, is to perform linear prediction in order to extend the time-series of correlation functions [79, 39, 80, 81]. While it is reported that this works remarkably well for simple spectra with few poles, we have not been able to successfully apply this to the correlation functions computed at non-zero temperatures in this work. For generic METTS states we mostly find strong disagreement between the actual time-evolution and the linear prediction. This is likely due to the fact that the correlators $\mathcal{C}_{i}(t)$ from METTS in general have a more complex structure than a superposition of a few sinusoidal waves. Thus, all the data we present is taken from actual time evolution. Recently, however, a more advanced method for performing linear prediction in the context of spectral functions with increased accuracy has been proposed [81]. It would be interesting to find out whether this method would be beneficial in the present context to extend the METTS time series.
The scaling of the method is determined by the maximal MPS bond dimension $\chi$ which needs to be simulated and the MPO bond dimension $D$ which typically remains finite. To perform both the METTS sampling as well as the time-evolution of METTS states, the proposed method has the same scaling as DMRG, which is $\mathcal{O}(\chi^{3}D+\chi^{2}D^{2})$, since we are using the TDVP time evolution method. The key novelty is, that due to the complex time evolution, the maximal necessary MPS bond dimension $\chi$ needed to resolve the time-evolved state is bounded, as opposed to real-time evolution, where the entanglement in general grows unbounded.
@@ Appendix A Projection error of TDVP
Figure 14: Imaginary time correlation function using TEBD and the 2-site TDVP approach, for $T/\Delta=0$ (a), $0.1$ (b) and $0.2$ (c). Two dashed lines, including $c_{2}e^{-\Delta\tau}$ ( $\Delta\approx 0.458$ ) and $c_{1}e^{-0.56\tau}$ are guides to the eye. The time evolution approach for measuring correlation function at $T=0$ is based on the ground state via DMRG calculation, where the ones at $T/J_{D}=0.1$ and $0.2$ are from averaging the METTS configurations.
One potential error of the TDVP method exists at the step of projecting the time-dependent Schrödinger equation onto the relevant MPS subspace [67]. This projection error is especially severe when time evolving an MPS with very low bond dimension. It turns out to be a serious issue for the Shastry-Sutherland model when the temperature is much lower than the first triplon gap $\Delta$. In this low-energy case, the MPS during the time evolution step does not differ significantly from the ground state, which is known to be a product state of spin singlets living on the off-diagonal bonds. The projection error leads to a biased estimation of the spin gap, even with the 2-site TDVP approach. On the other hand, time-evolving block decimation (TEBD) suffers only from the Trotter error and the MPS truncation error.
As a test, we show a comparison between TDVP and TEBD methods in Fig. 14. At low temperatures, one expects the imaginary time correlation function to decay with an exponential factor based on the estimated spin gap ($\Delta\approx 0.458J_{D}$) from the perturbation theory. However, as can be seen from the plot, $G(\tau)$ obtained from the 2-site TDVP method deviates from this expectation with a much steeper slope, for $T/\Delta=0$ and $0.1$. On the other hand, the long-time correlation from TEBD resolves the expected spin gap. Meanwhile, for temperatures comparable with the spin gap, enough entanglement is built up within the time-evolved MPS and we do not observe a severe projection error in TDVP. This can be seen from the consistent data between the two methods at $T/\Delta=0.2$. For the three temperatures here, we choose the time step in TEBD as $d\tau=0.1$ such that the Trotter error is sufficiently small.
@@ Appendix B Fictitious temperature dependence of stochastic analytic continuation
Appendix C Comparison of windowing functions
Figure 16: Comparison of spectral functions obtained via ED and the real-time METTS algorithm as in Fig. 4, but using a Parzen windowing function Eq. 91.
In this appendix we compare the results of computing the spectral functions using different windowing functions to perform the Fourier transform. In the main text in Fig. 4 we used the Hann window of the form,
$w(x)=0.5+0.5\cos\left(\pi x\right)\quad\text{for}\quad 0\leq|x|\leq 1.$ (90)
One could ask to what extent the result depends on the windowing function. To answer this question we compare two different windowing functions. First, we show results using the Parzen window of the form,
$w(x)=\begin{cases}1-6|x|^{2}+6|x|^{3}\quad\text{if}\quad 0\leq|x|\leq 1/2\\ 2(1-|x|^{3})\quad\text{if}\quad 1/2\leq|x|\leq 1\end{cases},$ (91)
in Fig. 16. The Parzen windowing function has previously been used in the context of time-dependent DMRG at zero temperature [20].
Figure 17: Comparison of spectral functions obtained via ED and the real-time METTS algorithm as in Fig. 4, but using a Gaussian windowing function Eq. 92 of width $\sigma=1/2$.
An equally valid choice of windowing function is a Gaussian window of the form,
$w(x)=\exp\left(-\frac{|x|^{2}}{2\sigma^{2}}\right)\quad\text{for}\quad 0\leq|x|\leq 1,$ (92)
where $\sigma$ controls the width of the Gaussian window. Results from applying the Gaussian windowing function with $\sigma=1/2$ are shown in Fig. 17.
Overall, we see that due to the chosen maximal simulation time $\Omega/J_{D}=50$ our results are not strongly affected by the choice of windowing function. The effects of the windowing function are expected to vanish for $\Omega\rightarrow\infty$.